OpenStudy (anonymous):

Algebraically determine the exact roots: 4x-1/x = 2x/3 help please? :( Midterm tomorrow!

OpenStudy (anonymous):

\[\frac{ 4x-1 }{ x }=\frac{ 2x }{ 3 }\] Cross multiply. \[3(4x-1)=2x(x)\] \[12x-3=2x^2\] Move everything to the RHS (Right Hand Side) \[2x^2-12x+3=0\] Solve the quadratic.

OpenStudy (anonymous):

Using the quadratic formula*

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

The root of the equation is when the graph of the function crosses the x axis. It is also where y = 0. So, just solve for x. 4x -1/x = 2x/3 4x^2 -1 = (2/3)(x^2) 4x^2 - (2/3)(x^2) - 1 = 0 12x^2 - 2x^2 -3 = 0 10 x^2 = 3 x^2 = 3/10 x = plus or minus sqrt(3/10)